Would you please help me with this question..
In triangle ABC, AC = 18, BC =10, and cosC = 1/2. Find the area of triangle ABC to the nearest tenth of a square unit.
Thanks!

Answer by stanbon(75887)

Would you please help me with this question..
In triangle ABC, AC = 18, BC =10, and cosC = 1/2. Find the area of triangle ABC to the nearest tenth of a square unit.
----------------
Assuming the triangle is a right triangle, since cos(C) = 1/2,
sin(C) = [sqrt(3)]/2
------------------------------
Formula:
Area = (1/2)ab*sin(C)
Area = (1/2)*10*18
= 5*9*sqrt(3)
= 77.94 sq. units
===================================
Cheers,
Stan H.